Please detailPDE Utt = Uzx + 2a sin(at) sin(1x) 0 < x <1 0<t< oo BCS S u(0,t) = 0 | u(1,t) = 0 0<t< oo ICs u(x,0) = 0 | u4(,0) = sin(nx) 0

Please detail
Answer #1

• Pelę U+-4an = 99 Sincort). Since) (0,+): 0 - UCLA) ; Octo U 42,0) - Sin 371 j05x (2,3) (4,5) ni Let Uc»= 2. an Ct). Sin (na. Sing U13,0) = 0 = O Sin (now) = anco) - Sininos) - On (0=0 · Also uiry,0) = S(X) - 2 enco) Sun(01029 = Žarco). Sın tnx) -